Last year Chloe quit her $60,000 per year job as a web-page designer for a leading computer software company to buy a small hotel on Saranac Lake. The purchase price of the hotel was $300,000, which she financed by selling a tax-free municipal bond that earned 5.5% per year Chloe’s total operating expenses and revenues were $100,000 and $200,000, respectively.

    a. Calculate Chloe’s accounting profit.

    b. Calculate Chloe’s economic profit.



1
Expert's answer
2021-09-12T19:37:18-0400

"Accounting Profit = Total Revenue - (Cost of Goods Sold + Operating Expenses + Taxes)"

= "\frac {5.5}{100} × 300000 = 16500"

"= 300000+ 16500+ (100000+200000)" "=" "316500-300000= 16500"


"Economic profit = 300000+ 16500-60000""= 256500"